OpenStudy (anonymous):

how do I work this out; find the equation of the straight line which passes through the point (0, 5) and is perpendicular to the line y=4x+5? I think it's y=5x-1 but I'm probably wrong...

OpenStudy (anonymous):

Using the knowlege of y=mx+b, where m is the slope and b is the y-intercept, the line y=4x+5 has a slope of 4. Perpendicular lines have negative reciprocal slopes, so the equation of the line you are looking for is y=-1/4 x + b. We still have to find the y-intercept. Luckily, we know the line passes through the point (0,5), so the y-intercept is 5. The final answer is y=-1/4x + 5.

OpenStudy (anonymous):

y=-1/4x+5